# See through the Dark: Learning Illumination-affined Representations for Nighttime Occupancy Prediction

### v2

@@ 1 Introduction
Understanding the 3D structure of the environment is a core task in autonomous driving [32, 48], as it allows vehicles to perceive their surroundings and make informed decisions. Recently, vision-based occupancy prediction [12, 16, 25, 26, 31, 39] has attracted growing interest due to its ability to estimate the spatial layout of occupied regions together with their semantic labels. Although existing methods perform well under daytime conditions, their performance drops significantly in nighttime scenes. The need for reliable perception in such low-light environments underscores the importance of advancing nighttime occupancy prediction. However, this task remains particularly challenging, as limited visibility and complex lighting conditions can severely degrade the quality of visual inputs.
As illustrated in Fig. 1(a), nighttime images not only exhibit low visibility but also contain both underexposed and overexposed regions. Underexposure due to insufficient illumination degrades visual features, whereas overexposure from artificial light sources, such as headlights and streetlamps, results in semantic deficiency in the saturated regions.
@@ 2 Related Work
Vision-based Occupancy Prediction. Recently, vision-based 3D occupancy prediction has attracted growing interest, with studies exploring both supervised and unsupervised learning paradigms [12, 15, 31, 33, 45]. Among supervised approaches, MonoScene [3] is a pioneering method for monocular occupancy prediction. BEVDet [11] adopts the Lift-Splat-Shoot (LSS) [35] for view transformation. BEVDet4D [9] further explores the temporal fusion strategy by fusing history frames. Building upon the BEVDet [11, 21, 22] series, FlashOcc [50] introduces a channel-to-height mechanism for memory-efficient occupancy prediction. In addition, several transformer-based methods have been proposed. For instance, BEVFormer [25] utilizes spatiotemporal transformers to construct BEV features. SparseOcc [27] takes the first step to explore the fully sparse architecture. For unsupervised methods [13, 52], SelfOcc [13] and OccNeRF [52] are two representative methods that employ volume rendering to generate self-supervised signals. Recently, Gaussian-based representations [6, 15, 42] have emerged as a powerful approach for 3D scene modeling. For example, GaussianTR [15] aligns Gaussian features with foundation models, achieving state-of-the-art zero-shot performance.
@@
Low-light Image Enhancement. Low-light image enhancement [8, 37, 41, 49] aims to improve the brightness and contrast. Histogram Equalization (HE) [36], along with its variant CLHE [36], is a classical method for enhancing the global contrast of images. However, their effectiveness diminishes in the presence of significant image noise or non-uniform luminance. Retinex-based methods [37] offer an alternative strategy. Retinex theory [18] assumes that the low-light image can be decomposed into illumination and reflectance, where the reflectance is regarded as the enhanced image. Recently, network-based Retinex methods [2, 28, 37, 43] combine convolutional neural networks (CNNs) with Retinex theory to achieve improved accuracy. However, in real-world nighttime driving scenarios, the presence of artificial light sources often results in spatially uneven illumination. In such cases, indiscriminate enhancement may degrade the visual quality of already well-lit images.
@@
View Transformation. Transforming 2D image features into 3D space is a crucial step in 3D perception tasks. Bottom-up methods [21, 22, 35, 45] rely on depth estimation to project multi-view features into the BEV space, a paradigm first introduced by LSS [35]. However, the resulting BEV features are typically sparse, posing challenges for tasks that require dense spatial representations. In contrast, top-down transformer-based methods [19, 23, 25, 44] generate dense BEV features, but their lack of explicit geometric constraints makes them vulnerable to occlusions and depth inconsistencies. Recent studies [20, 24, 31, 38, 51] attempt to integrate the strengths of both paradigms. For instance, FB-OCC [24] leverages LSS-generated features to initialize transformer queries, while COTR [31] extends this strategy by employing compact occupancy representations to preserve rich geometric information. Nevertheless, these methods primarily address view transformation from a projection-based perspective, neglecting the influence of environmental factors such as illumination.
@@ 3.2 Selective Low-light Image Enhancement
[…]
@@ 4 Experiments
Dataset. We evaluate our method on both real and synthetic nighttime scenarios. (1) Occ3d-nuScenes [1, 39] includes 700 scenes for training and 150 for validation, with annotations spanning a spatial range of -40m to 40m along both the X and Y axes, and -1m to 5.4m along the Z axis. The occupancy labels are defined using voxels with a size of $0.4\rm m\times 0.4\rm m\times 0.4\rm m$, containing 17 distinct categories. Notably, the training and validation sets include 84 and 15 real-world nighttime scenes, respectively. (2) nuScenes-C [46, 47] is a synthetic benchmark that introduces eight types of data corruptions, each applied at three intensity levels to the validation set of nuScenes. Among them, the Dark corruption simulates nighttime by reducing overall brightness and contrast and adding random noise.
@@ 5 Conclusion
Limitation. While LIAR achieves promising results in nighttime scenarios, it relies on the accuracy of illumination map estimation, which may not be accurate under extremely irregular lighting conditions. Moreover, current occupancy prediction datasets contain only a limited number of nighttime data, constraining comprehensive evaluation of model performance in such settings. Future work is needed to assess the generalizability of the model across more diverse and challenging environments.
@@ Appendix A Technical Appendices and Supplementary Material
       Method        Flops (G)        Training Memory (M)        mIoU
       BEVDetOcc (2f) [9]        541.21        4828        15.86
       FlashOcc (2f) [50]        439.71        2834        18.15
       COTR (2f) [31]        740.89        12439        20.01
       Our LIAR (2f)        690.40        5177        22.09
@@ A.1 Metrics
A.2 Comparison of SLLIE and Other Methods
@@
A.3 Computational Cost
A.4 Additional Visual Results
Fig. 10 shows visual comparisons on the same scene under all three severity levels on the nuScenes-C, while Fig. 11 illustrates results on real-world nighttime data from the Occ3D-nuScenes.

### v3

@@ 1 Introduction
Understanding the 3D structure of the environment [15, 16, 52, 53, 54, 55] is a core task in autonomous driving, as it allows vehicles to perceive their surroundings and make informed decisions. Recently, vision-based occupancy prediction [12, 18, 26, 31] has attracted growing interest due to its ability to estimate the spatial layout of occupied regions together with their semantic labels. Although existing methods perform well under daytime conditions, their performance drops significantly in nighttime scenes. The need for reliable perception in such low-light environments underscores the importance of advancing nighttime occupancy prediction. However, this task remains particularly challenging, as limited visibility and complex lighting conditions can severely degrade the quality of visual inputs.
As illustrated in Fig. 1(a), nighttime images not only exhibit low visibility but also contain both underexposed and overexposed regions [23, 39, 57]. Underexposure caused by insufficient illumination severely degrades visual features, whereas overexposure from artificial light sources, such as vehicle headlights and streetlamps, results in significant semantic deficiency in the saturated regions.
@@ 2 Related Work
Vision-based Occupancy Prediction. Recently, vision-based 3D occupancy prediction has attracted growing interest, with studies exploring both supervised and unsupervised learning paradigms [12, 31, 32, 49, 64, 68]. Among supervised approaches, MonoScene [3] is a pioneering method for monocular occupancy prediction. BEVDet [11] adopts the Lift-Splat-Shoot (LSS) [34] for view transformation. BEVDet4D [9] further explores the temporal fusion strategy by fusing history frames. Building upon the BEVDet [11, 20, 21] series, FlashOcc [63] introduces a channel-to-height mechanism for memory-efficient occupancy prediction. In addition, several transformer-based methods have been proposed. For instance, BEVFormer [25] utilizes spatiotemporal transformers to construct BEV features. SparseOcc [27] takes the first step to explore the fully sparse architecture. For unsupervised methods [13, 66], SelfOcc [13] and OccNeRF [66] are two representative methods that employ volume rendering to generate self-supervised signals. Recently, Gaussian-based representations [6, 17, 42] have emerged as a powerful approach for 3D scene modeling. For example, GaussianTR [17] aligns Gaussian features with foundation models, achieving state-of-the-art zero-shot performance.
@@
Low-light Image Enhancement. Low-light image enhancement [8, 23, 36, 41, 48, 57] aims to improve the brightness and contrast. Histogram Equalization [35], along with its variant CLHE [35], is a classical method for enhancing the global contrast of images. However, their effectiveness diminishes in the presence of significant image noise or non-uniform illumination. Retinex-based methods [36] offer an alternative strategy. Recently, network-based Retinex methods [2, 28, 36, 46] combine convolutional neural networks with Retinex theory to achieve improved accuracy. Nevertheless, in real-world nighttime driving scenarios, artificial light sources often cause spatially uneven illumination [23, 39]. To address this issue, LCDPNet [39] leverages local color distributions to correct illumination imbalance, while CSEC [23] models color distribution shifts to enhance image quality. Different from the aforementioned methods, we do not uniformly treat all nighttime images as requiring enhancement. Instead, we selectively enhance images that suffer from poor illumination, as indiscriminate enhancement may degrade the visual quality of already well-lit ones.
@@
View Transformation. Transforming 2D image features into 3D space is a crucial step in 3D perception tasks [58, 59, 60, 61, 62]. Bottom-up methods [20, 21, 34] rely on depth estimation [43, 44, 45, 56] to project multi-view features into the BEV space, a paradigm first introduced by LSS [34]. However, the resulting BEV features are typically sparse, posing challenges for tasks that require dense spatial representations. In contrast, top-down transformer-based methods [22, 25, 47] generate dense BEV features, but their lack of explicit geometric constraints makes them vulnerable to occlusions and depth inconsistencies. Recent studies [24, 31, 37, 65] attempt to integrate the strengths of both paradigms. For instance, FB-OCC [24] leverages LSS-generated features to initialize transformer queries, while COTR [31] extends this strategy by employing compact occupancy representations to preserve rich geometric information. Nevertheless, these methods primarily address view transformation from a projection-based perspective, neglecting the influence of environmental factors such as illumination.
@@ 3.2 Selective Low-light Image Enhancement
where $\lfloor\cdot\rfloor$ denotes the floor operation to discretize continuous coordinates into pixel indices. This aggregated $\boldsymbol{S}$ provides an informative estimation of 3D illumination distribution, enabling the model to localize overexposed regions and emphasize them during subsequent feature reasoning.
@@ 4 Experiments
Dataset. We evaluate our method on both real and synthetic nighttime scenarios. (1) Occ3D-nuScenes [38] includes 700 scenes for training and 150 for validation, with annotations spanning a spatial range of -40m to 40m along both the X and Y axes, and -1m to 5.4m along the Z axis. The occupancy labels are defined using voxels with a size of $0.4\rm m\times 0.4\rm m\times 0.4\rm m$, containing 17 categories. Notably, the training and validation sets include 84 and 15 real-world nighttime scenes, respectively. (2) nuScenes-C [51] is a synthetic benchmark that introduces eight types of data corruptions, each applied at three intensity levels to the validation set of nuScenes. Among them, the Dark corruption simulates nighttime by reducing brightness and contrast and adding random noise.
@@ 5 Conclusion
Limitation. We evaluate our method on both real and synthetic nighttime scenarios using Occ3D-nuScenes [38] and nuScenes-C [50], respectively. For Occ3D-nuScenes, although it serves as the most widely adopted benchmark for occupancy prediction, its nighttime subset is relatively small, and several categories are absent. These limitations may reduce the diversity of evaluated scenarios. As for nuScenes-C, it is a synthetic nighttime dataset generated by darkening daytime images and injecting noise. Due to its synthetic nature, nuScenes-C inevitably differs from real-world data. This domain gap can be broadly categorized into two aspects: First, it retains the semantics and motion patterns of daytime scenes, which differ from actual nighttime driving scenarios. Second, it fails to replicate sensor-induced degradations, such as thermal noise, motion blur, and compression artifacts.
@@
6 Acknowledgment
This work was supported by the National Science Fund of China under Grant Nos. U24A20330, 62361166670, and 62306155, and by the Technology Major Project No. 2022ZD0116305.
@@ Appendix A Technical Appendices and Supplementary Material
Method Flops (G) Memory (GB) Params (M) FPS (img/s) mIoU
BEVDetOcc [9] 541.21 4.71 34.97 1.1 21.98
FlashOcc [63] 439.71 2.76 58.67 7.5 23.40
OPUS [40] 215.54 2.00 73.17 22.4 20.28
COTR [31] 740.89 12.12 37.71 0.5 25.17
Our LIAR (2f) 690.40 5.05 64.39 4.7 27.33
@@ A.1 Metrics
A.2 Computational Cost
A.3 SLLIE Pretraining
The SLLIE module is pretrained on the nighttime split of the nuScenes training set in a self-supervised manner. Following SCI [30], we use the fidelity loss as follows:
$\mathcal{L}_{f}=\sum_{t=1}^{T}\left\|\mathrm{x}^{t}-\left(\mathrm{y}+\mathrm{s}^{t-1}\right)\right\|^{2},$ (16)
where $\mathrm{y}$ is the input low-light image, $\mathrm{x}^{t}$ is the is the illumination estimate at stage $t$ and $\mathrm{s}^{t-1}$ is the output of the self-calibrated module from the previous stage. When the input image $\mathrm{y}$ is already well-exposed (e.g., daytime image), the pseudo-target $\mathrm{y}+\mathrm{s}^{t-1}$ becomes over-bright, thereby misleading the network to over-enhance the illumination $\mathrm{x}^{t}$. Consequently, $\mathcal{L}_{f}$ generates erroneous gradients that hinder the model’s ability to generalize to genuine low-light scenarios. Given that the full training set comprises a large proportion of daytime images (approximately 88%), jointly training the SLLIE module with the rest of the occupancy model results in optimization collapse. Therefore, we pretrain the SLLIE module on nighttime data and freeze its weights during training to ensure stability.
A.4 Comparison of SLLIE and Other Methods
@@
Moreover, we conduct experiments where we apply state-of-the-art low-light enhancement methods as a preprocessing step before standard occupancy models. Specifically, we use LCDPNet [39] and CSEC [23] as low-light enhancement modules. As shown in Tab. 7, while applying enhancement methods leads to marginal improvements, our LIAR consistently outperforms existing methods. For example, our LIAR outperforms COTR + LCSDPNet and COTR + CSEC by 1.87 and 1.92 mIoU, respectively. This highlights that the improvements are not solely attributable to low-light preprocessing, but also stem from our illumination-aware designs.
Table 6: Comparison with low-light enhancement methods on the Occ3D-nuScenes.
Method mIoU
BEVDetOcc [9] 15.86
BEVDetOcc + LCDPNet [39] 16.18 (+0.32)
BEVDetOcc + CSEC [23] 16.23 (+0.37)
FlashOcc [63] 18.15
FlashOcc + LCDPNet 18.29 (+0.14)
FlashOcc + CSEC 18.23 (+0.08)
COTR [31] 20.01
COTR + LCDPNet 20.22 (+0.21)
COTR + CSEC 20.17 (+0.16)
Our LIAR 22.09
Table 7: Performance on the Occ3D-nuScenes.
Method mIoU
all day night
BEVDetOcc (2f) [9] 36.10 36.85 21.98
BEVFormer (4f) [25] 23.67 24.17 13.77
FlashOcc (2f) [63] 37.84 38.90 23.40
OPUS (8f) [40] 33.20 33.96 20.28
SparseOcc (8f) [27] 31.08 31.59 22.79
COTR (2f) [31] 38.70 39.47 25.17
Our LIAR (2f) 39.57 40.42 27.33
A.5 Performance on Daytime Scenes
While our method is primarily designed for nighttime scenarios, we also evaluate our method on daytime scenes. Tab. 7 shows that our LIAR consistently outperforms existing methods on both the daytime and full datasets, surpassing the second-best COTR [31] by 0.87 and 0.95 mIoU, respectively. These results highlight the generalizability of our approach beyond nighttime conditions.
A.6 Additional Visual Results
Fig. 10 illustrates results on real-world nighttime data from the Occ3D-nuScenes, while Fig. 11 shows visual comparisons on the same scene under all three severity levels on the nuScenes-C.